# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 10.3. A purely topological characterization of filters
As we discussed in Sections 2.3 and 2.4, the combinatorial structure of filters imposes strong constraints on their topology. But is it possible to go the other way? In other words, are there natural conditions on a space $X$ which will guarantee that $X$ is homeomorphic to a filter? Van Engelen showed that, in the Borel context, this question has a very elegant positive answer (see [vE5, Theorem 3.4]). The following result shows that, under suitable determinacy assumptions, his characterization extends beyond the Borel realm.
@@
The implication $(1)\rightarrow(2)$ follows from the discussion in Section 2.3, Theorem 2.10, and Corollary 2.12. In order to prove the implication $(2)\rightarrow(1)$, assume that condition $(2)$ holds. If $X$ is a $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$ space, then the desired result follows from [vE5, Lemma 3.3], so assume that this is not the case. By Theorem 9.7, there exists a good Wadge class $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$ such that $X$ has exact complexity $\mathbf{\Gamma}$. By Theorems 9.18 and 10.4, it will be enough to show that $\mathbf{\Gamma}$ is closed under squares. To see this, proceed as in the proof of Theorem 10.4. ∎
@@ 10.4. Filters versus topological groups
As Theorem 9.19 shows, the lower-right portion of the diagram from Section 2.4 “collapses” under $\mathsf{AD}$. Therefore, it seems natural to ask whether the rest of the diagram could collapse as well. More precisely, one might ask: does $\mathsf{AD}$ imply that every zero-dimensional topological group is homemorphic to a filter?
As usual, the trivial case of locally compact spaces provides several counterexamples (see the spaces mentioned by Proposition 2.14). And even after excluding these spaces, the question remains too naive, as $\omega^{\omega}$ is a zero-dimensional topological group that is not homeomorphic to a filter (by Corollary 2.12). However, adding the requirement that the topological group is meager yields an open question. In fact, even the Borel version of this problem is still open. It was first formulated at the very end of [vE6], where van Engelen solved the problem for spaces of low complexity (see [vE6, Theorem 3.4]). The precise statements are as follows.111 We added the assumption “not locally compact” in the statement of Theorem 10.6 because our definition of filter is slightly more restrictive than that of van Engelen. However, this does not result in any loss of generality (see [MZ1, Proposition 2]).
@@
However, Conjecture 10.8 would be false if we altogether dropped the determinacy assumption. In fact, working in $\mathsf{ZFC}$, it is a simple exercise to construct an uncountable zero-dimensional topological group $X$ in which $2^{\omega}$ does not embed. It is easy to realize that the space ×$\mathbb{Q}\times X$ will be a meager topological group that is not homeomorphic to a filter.
We conclude this section by showing that, although the diagram from Section 2.4 might partially collapse, under no circumstance will a complete collapse happen. We will combine ideas from [Me4, Proposition 5.4] and [vE3, Theorem 3.6.5].
@@
We remark that the semifilter $\mathcal{T}$ described in the above proof is homeomorphic to the notable space $\mathbf{T}$ introduced by van Douwen (see [vEvM] and [Me4, Proposition 5.4]). Since ×$\mathbb{Q}\times\mathbf{T}$ is not a topological group by [vE4, Corollary 4.3], and multiplication by $\mathbb{Q}$ preserves the property of being homeomorphic to a semifilter by [Me4, Lemma 8.2], this shows that that the example given by Proposition 10.10 could be made meager as well.
@@
The following is a list of most of the symbols and terminology used in this monograph, organized by the section in which they are defined.
  • Section 1.1:
@@
  • Section 2.1:
    power-set $\mathcal{P}(\Omega)$, $\mathsf{Fin}(\Omega)$, $\mathsf{Cof}(\Omega)$, $\Omega^{<\omega}$, $\Omega^{\leq\omega}$, $\mathsf{N}_{s}$, concatenation $s^{\frown}t$, monotone, eventually growing, induced function $f_{\sigma}$, image $f[S]$, inverse image $f^{-1}[S]$, domain $\mathsf{dom}(f)$, range $\mathsf{ran}(f)$, identity function $\mathsf{id}_{X}$, retraction, clopen, base, zero-dimensional, codense, meager subset, comeager subset, meager space, Baire space, embedding, dense embedding, closed function, open function, open ball $\mathsf{B}(z,\varepsilon)$, diameter $\mathsf{diam}(S)$, mesh $\mathsf{mesh}(\mathcal{S})$, compatible metric, $\mathsf{Borel}(Z)$, Borel space, $\mathbf{\Sigma}^{0}_{\xi}$-measurable function, Borel function, differences operation $\mathsf{D}_{\eta}$, class of differences $\mathsf{D}_{\eta}(\mathbf{\Sigma}^{0}_{\xi}(Z))$.
  • Section 2.2:
@@
  • Section 2.4:
    strongly homogeneous, local base, $\pi$-base.
  • Section 3.1:
    $\widecheck{\mathbf{\Gamma}}$, selfdual class, $\Delta(\mathbf{\Gamma})$, Wadge-reduction $\leq$, strict Wadge-reduction $<$, selfdual set, Wadge class $A\!\!\downarrow\,$, continuously closed, $\mathsf{NSD}(Z)$, $\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$.
  • Section 3.2:
@@
  • Section 3.5:
@@
  • Section 4.1:
    Hausdorff operation $\mathcal{H}_{D}(A_{0},A_{1},\ldots)$, Hausdorff class $\mathbf{\Gamma}_{D}(Z)$.
  • Section 4.2:
    relativizable pointclass, induced relativizable pointclass, vertical section $S_{z}$, $W$-universal set.
  • Section 5.1:
    (first) separation property, second separation property.
  • Section 6.1:
@@
  • Section 8.1:
    Knaster-Reichbach cover ($\mathsf{KR}$-cover).
  • Section 8.2:
    Knaster-Reichbach system ($\mathsf{KR}$-system).
  • Section 8.3:
    reasonably closed Wadge class.
  • Section 9.1:

### v3

@@ 7.3. A purely topological characterization of filters
As we discussed in §2.3 and §2.4, the combinatorial structure of filters imposes strong constraints on their topology. But is it possible to go the other way? In other words, are there natural conditions on a space $X$ which will guarantee that $X$ is homeomorphic to a filter? Van Engelen showed that, in the Borel context, this question has a very elegant positive answer (see [vE5, Theorem 3.4]). The following result shows that, under suitable determinacy assumptions, his characterization extends beyond the Borel realm.
@@
The implication $(1)\rightarrow(2)$ follows from the discussion in §2.3, Theorem 2.9, and Corollary 2.11. In order to prove the implication $(2)\rightarrow(1)$, assume that condition $(2)$ holds. If $X$ is a $\Delta(\mathsf{D}_{\omega}(\mathbf{\Sigma}^{0}_{2}))$ space, then the desired result follows from [vE5, Lemma 3.3], so assume that this is not the case. By Theorem 6.7, there exists a good Wadge class $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$ such that $X$ has exact complexity $\mathbf{\Gamma}$. By Theorems 6.17 and 7.4, it will be enough to show that $\mathbf{\Gamma}$ is closed under squares. To see this, proceed as in the proof of Theorem 7.4. ∎
@@ 7.4. Filters versus topological groups
As Theorem 6.18 shows, the lower-right portion of the diagram from §2.4 “collapses” under $\mathsf{AD}$. Therefore, it seems natural to ask whether the rest of the diagram could collapse as well. More precisely, one might ask: does $\mathsf{AD}$ imply that every zero-dimensional topological group is homemorphic to a filter?
As usual, the trivial case of locally compact spaces provides several counterexamples (see the spaces mentioned by Proposition 2.13). And even after excluding these spaces, the question remains too naive, as $\omega^{\omega}$ is a zero-dimensional topological group that is not homeomorphic to a filter (by Corollary 2.11). However, adding the requirement that the topological group is meager yields an open question. In fact, even the Borel version of this problem is still open. It was first formulated at the very end of [vE6], where van Engelen solved the problem for spaces of low complexity (see [vE6, Theorem 3.4]). The precise statements are as follows.191919 We added the assumption “not locally compact” in the statement of Theorem 7.6 because our definition of filter is slightly more restrictive than that of van Engelen. However, this does not result in any loss of generality (see [MZ1, Proposition 2]).
@@
However, Conjecture 7.8 would be false if we altogether dropped the determinacy assumption. In fact, working in $\mathsf{ZFC}$, it is a simple exercise to construct an uncountable zero-dimensional topological group $X$ in which $2^{\omega}$ does not embed. It is easy to realize that the space $\mathbb{Q}\times X$ will be a meager topological group that is not homeomorphic to a filter.
We conclude this subsection by showing that, although the diagram from §2.4 might partially collapse, under no circumstance will a complete collapse happen. We will combine ideas from [Me4, Proposition 5.4] and [vE3, Theorem 3.6.5].
@@
We remark that the semifilter $\mathcal{T}$ described in the above proof is homeomorphic to the notable space $\mathbf{T}$ introduced by van Douwen (see [vEvM] and [Me4, Proposition 5.4]). Since $\mathbb{Q}\times\mathbf{T}$ is not a topological group by [vE4, Corollary 4.3], and multiplication by $\mathbb{Q}$ preserves the property of being homeomorphic to a semifilter by [Me4, Lemma 8.2], this shows that that the example given by Proposition 7.10 could be made meager as well.
@@
The following is a list of most of the symbols and terminology used in this article, organized by the subsection in which they are defined.
  • §1.1:
@@
  • §2.1:
    power-set $\mathcal{P}(\Omega)$, $\mathsf{Fin}(\Omega)$, $\mathsf{Cof}(\Omega)$, $\Omega^{<\omega}$, $\mathsf{N}_{s}$, monotone, image $f[S]$, inverse image $f^{-1}[S]$, identity function $\mathsf{id}_{X}$, retraction, base, local base, $\pi$-base, clopen, zero-dimensional, meager subset, comeager subset, meager space, Baire space, embedding, dense embedding, closed function, open function, $\mathsf{Borel}(Z)$, Borel space, $\mathbf{\Sigma}^{0}_{\xi}$-measurable function, Borel function, differences operation $\mathsf{D}_{\eta}$, class of differences $\mathsf{D}_{\eta}(\mathbf{\Sigma}^{0}_{\xi}(Z))$.
  • §2.2:
@@
  • §2.4:
    strongly homogeneous.
  • §3.1:
    $\widecheck{\mathbf{\Gamma}}$, selfdual class, $\Delta(\mathbf{\Gamma})$, Wadge-reduction $\leq$, strict Wadge-reduction <<, selfdual set, Wadge class $A\!\!\downarrow\,$, continuously closed, $\mathsf{NSD}(Z)$, $\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$.
  • §3.2:
@@
  • §3.5:
    Hausdorff operation $\mathcal{H}_{D}(A_{0},A_{1},\ldots)$, Hausdorff class $\mathbf{\Gamma}_{D}(Z)$, $W$-universal set, vertical section $S_{y}$, relativizable pointclass.
  • §3.7:
    expanded Hausdorff class $\mathbf{\Gamma}^{(\xi)}_{D}(Z)$.
  • §3.8:
    (first) separation property, second separation property.
  • §3.9:
@@
  • §4.1:
@@
  • §6.1: